OpenStudy (anonymous):

If g(x) = x2 + 3, find g(4). A. 11 B. 19 C. 16 D. 8

OpenStudy (jacob902):

b

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle g(x)=x^2+3 }\) if I wanted to find \(g(5)\) i would simply plug in 5 for x. This way: \(\large\color{black}{ \displaystyle g(\color{red}{5})=\color{red}{5}^2+3 }\) \(\large\color{black}{ \displaystyle g(5)=25+3 }\) \(\large\color{black}{ \displaystyle g(5)=28 }\)

OpenStudy (jacob902):

Way of getting answer is substitute the values your are given in position of x and then calculate. Example; 1) g(x)=x2+3

OpenStudy (anonymous):

thnx :)

OpenStudy (solomonzelman):

TO find g(4), all you got to do is to plug in 4 instead of x, just the same way as i did with g(5).

OpenStudy (anonymous):

i got it thnx :)